OpenStudy (maheshmeghwal9):

If x^3+ax+b=0, & a,b belongs to real numbers, b is not equal to zero.Then, why a<0,since it is the a statement which holds true.And all the roots of this equation are real.

OpenStudy (blockcolder):

So your question is...?

OpenStudy (maheshmeghwal9):

my ques is why a<0.under what coditions?

OpenStudy (maheshmeghwal9):

sorry,under what conditions?

OpenStudy (maheshmeghwal9):

this is a question of theory of equations.

OpenStudy (blockcolder):

Do you have other assumptions on the cubic equation, like, how many solutions does it have, or other stuff like that?

OpenStudy (maheshmeghwal9):

no I don't know about the solutions but I know that this is a thoughtful question.

OpenStudy (maheshmeghwal9):

this is only verify and prove question.

OpenStudy (anonymous):

The statement is not complete. For example: we say the standard form of quadratic equation is \[ax^2+bx+c=0, where a, b, c \in R \] and a, b not both zero

OpenStudy (maheshmeghwal9):

but question is is true because my sir has solved it and i have lost that paper on which sir solved that.

OpenStudy (freckles):

I don't understand the question. You want us to solve for x? Is that what your asking?

OpenStudy (maheshmeghwal9):

what is this?

OpenStudy (maheshmeghwal9):

one more thing about this quetion

OpenStudy (maheshmeghwal9):

all the roots of this cubic equations are real

OpenStudy (freckles):

Oh ok now the question makes more sense. lol/

OpenStudy (freckles):

Omg I'm sorry. My page got killed when I was typing all of that. :(

OpenStudy (maheshmeghwal9):

np!

OpenStudy (freckles):

@foolformath what do you think of this problem? For some reason I was thinking about finding f' And I found critical numbers pm sqrt(-a/3) I found it is increasing on (-inf,-sqrt(-a/3)) and (sqrt(-a/3),inf) and decreasing on (-sqrt(-a/3),sqrt(-a/3)) I can't decide what to do from this or if i can do anything with what i found....

OpenStudy (maheshmeghwal9):

ok np. :(

OpenStudy (freckles):

i'm still thinking :(

OpenStudy (freckles):

@KingGeorge any thoughts?

OpenStudy (kinggeorge):

Sorry for asking this, but just to clarify, you're asking for conditions under which a will be less than 0 correct?

OpenStudy (freckles):

What do you think about taking the contrapositive?

OpenStudy (kinggeorge):

If that's the question, looking at the derivative is a good idea. \[f(x)=x^3+ax+b\]\[f'(x)=3x^2+a\]If \(a\) is negative, then we would have 2 critical points that translate to local maxima and local minima. If a is positive, we would have 0 critical points.

OpenStudy (kinggeorge):

If we drew the graphs, \(a\) negative would look similar to this:And \(a\) positive would look someting like